OpenStudy (anonymous):

Solving a system? 16x^2 - 3y^2 = -11 8x - y = -11 The answer for above are: (-1,3) and (-2,-5) but idk how they got that 2. x^2 +9y -10x +36y = 20 x- 3y = 2

OpenStudy (anonymous):

Use substitution.

OpenStudy (anonymous):

16x^2 - 3y^2 = -11 8x - y = -11 16x^2 - 3y^2 = -11 8x - y = -11 -> -y = -8x - 11 -> y = 8x + 11 16x^2 - 3y^2 = -11 y = 8x + 11 16x^2 - 3(8x + 11)^2 = -11 Continue.

OpenStudy (anonymous):

i did do that but i didnt get the correct answer.

OpenStudy (anonymous):

16x^2 - 3(8x + 11)^2 = -11 16x^2 - 3(8x + 11)^2 + 11 = 0

OpenStudy (anonymous):

Did you do that also? @Srini143

OpenStudy (anonymous):

yep!

OpenStudy (anonymous):

thats where i get stuck

OpenStudy (anonymous):

16x^2 - 3(8x + 11)^2 = -11 16x^2 - 3(8x + 11)^2 + 11 = 0 16x^2 - 3(8x + 11)(8x + 11) + 11 = 0 16x^2 + (-3(64x^2 + 176x + 121))+11 = 0 16x^2 + (-192x^2 - 528x - 363) + 11 = 0 -176x^2 - 528x - 363 + 11 = 0 -176x^2 - 528x - 352 = 0 Are you following?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

but i solved after that and i got this super long and complex answer

OpenStudy (anonymous):

What did you get?

OpenStudy (anonymous):

\[\left\{8 x-y=-11, y\text{=}11+8 x, y^2=64 x^2+176 x+121\right\}\]\[16 x^2-3\left(64 x^2+176 x+121\right)\text{=}-11\]\[-363-528 x-176 x^2=-11\]\[-176 x^2-528 x-352=0 \]The next surprising result was discovered using Mathematica.\[-176 (1+x) (2+x)=0\]\[{x=-1 , x=-2} \]

OpenStudy (anonymous):

wait whattt

OpenStudy (anonymous):

^ yup like Rob said, you get to x values. that is what i got too.

OpenStudy (anonymous):

Thank u sooo much!!!

OpenStudy (anonymous):

X = -2 X = -1

OpenStudy (anonymous):

Now you plug in and substitute those x values back into one of the equations and solve for y, to get the y values.

OpenStudy (anonymous):

{Best Response} ?